If the coordinates of the points A and B are (2,4) and (4,2) respectively and point M in such that A-M-B also AB=3AM then the coordinates of M are

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

points A and B are (2,4) and (4,2)

given

AB=3AM

AB/3 =AM/1

A bisect the point B and M externally.

let coordinate of point M is (x,y)

2 = (3*x -1*4)/(3-1)         and    4 = (3*y -1*2)/(3-1)

x=8/3                                    and y  =10/3

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin